OpenStudy (anonymous):

If iron ore contains 20% of the iron oxide (Fe2O3) - How many tons of it is required to produce one ton of iron

OpenStudy (aaronq):

Find the mass percent composition of Iron in the ore. -> mass % (Fe in Fe2O3) http://www.chem.tamu.edu/class/majors/tutorialnotefiles/percentcomp.htm ------------------- Since 20% of m(ore) is Fe2O3 = m(Fe2O3) = m(ore)*0.2 mass % (Fe in Fe2O3)*[m(ore)*0.2]=1 ton=1000 kg I'll re-write it \[masspercent_{Fe}*(m _{ore}*0.2) = 1000kg =1,000,000 g\]

OpenStudy (aaronq):

you're solving for m(ore)